# Small-Signal Stability Oriented Real-Time Operation of Power Systems with a High Penetration of Inverter-Based Resources

### v1

@@ III-C Regression-based Stability Constraint Training
The training dataset is generated using the methodology proposed in [rossiML2022], by varying demand and generation within the operating ranges specified in Tables III. The stability assessment of each operating point is executed by using the eigenvalues-based analysis provided by the STAMP tool [arevalo2025matlab]. Figure 2 presents a 3D scatter plot of the dataset instances. On the left, the samples are shown as a function of the total power demand ($P_{D}$), the share of this demand supplied by the SG, and the distribution of the remaining demand between the GFM and GFL converters. Analysis of these instances indicates that system loading above 400 MW tends to induce instability. Increasing the share of power supplied by the SG mitigates this effect, while, for equal SG participation, a lower contribution from the GFM converter enhances stability. On the right, the training samples are represented as a function of the voltage setpoints of the three generators. This visualization reveals that when the voltage of $\text{GFM}_{1}$ exceeds 1 p.u., the system becomes unstable. Given this insight into the system’s dynamic behavior, an OPF is formulated in this study with a voltage constraint at bus 1 of $0.9\leq V_{1}\leq 1.0$ p.u., and its solutions are compared with those obtained from the SSSC-OFO and SSSC-OPF formulations, which incorporate the regression-based stability constraint. However, it is important to note that such relationships are relatively easy to identify in this small 9-bus system, whereas in larger networks this approach would be impractical, as these dependencies become much less apparent.
@@
Before training, the data are preprocessed by applying feature selection, retaining as model inputs only uncorrelated variables. Model accuracy is then evaluated using an 80/20 split of the dataset for training and testing, respectively. The resulting $R^{2}$ score is 0.999 which confirms the suitability of the regression model for this task. Finally, the model is trained using the full dataset. The training process follows the two-step MARS fitting procedure: a forward selection phase and a backward pruning phase [hastie2009elements]. The pruning step reveals that only two quantities are sufficient to fit a model that accurately computes the DI, namely the voltage magnitudes at buses 1 and 6. The resulting expression is therefore:
@@ III-D SSSC-OFO Validation
Medium
Demand Case
($P_{D}=318.55$ MW)
OPF 239.54 54 25 10.45 -16.64 -32.08 0.6589 1.0014 1.0015
@@
High
Demand Case
($P_{D}=661.41$ MW)
OPF 486.40 56.32 118.70 161.28 47.73 35.90 0.2239 1.0020 1.0020

### v2

@@ III-C Regression-based Stability Constraint Training
The training dataset is generated using the methodology proposed in [30], by varying demand and generation within the operating ranges specified in Tables III. The stability assessment of each operating point is executed by using the eigenvalues-based analysis provided by the STAMP tool [2]. Figure 2 presents a 3D scatter plot of the dataset instances. On the left, the samples are shown as a function of the total power demand ($P_{D}$), the share of this demand supplied by the SG, and the distribution of the remaining demand between the GFM and GFL converters. Analysis of these instances indicates that system loading above 400 MW tends to induce instability. Increasing the share of power supplied by the SG mitigates this effect, while, for equal SG participation, a lower contribution from the GFM converter enhances stability. On the right, the training samples are represented as a function of the voltage setpoints of the three generators. This visualization reveals that when the voltage of $\text{GFM}_{1}$ exceeds 1 p.u., the system becomes unstable. Given this insight into the system’s dynamic behavior, an OPF is formulated in this study with a voltage constraint at bus 1 of $0.9\leq V_{1}\leq 1.0$ p.u., and its solutions are compared with those obtained from the SSSC-OFO and SSSC-OPF formulations, which incorporate the regression-based stability constraint. However, it is important to note that such relationships are relatively easy to identify in this small 9-bus system, whereas in larger networks this approach would be impractical, as these dependencies become much less apparent.
@@
Before training, the data are preprocessed by applying feature selection, retaining as model inputs only uncorrelated variables. Model accuracy is then evaluated using an 80/20 split of the dataset for training and testing, respectively. The resulting $R^{2}$ score is 0.999 which confirms the suitability of the regression model for this task. Finally, the model is trained using the full dataset. The training process follows the two-step MARS fitting procedure: a forward selection phase and a backward pruning phase [15]. The pruning step reveals that only two quantities are sufficient to fit a model that accurately computes the DI, namely the voltage magnitudes at buses 1 and 6. The resulting expression is therefore:
@@ III-D SSSC-OFO Validation
Medium Demand Case ($P_{D}=318.55$ MW) OPF 239.54 54 25 10.45 -16.64 -32.08 0.6589 1.0014 1.0015
@@
High Demand Case ($P_{D}=661.41$ MW) OPF 486.40 56.32 118.70 161.28 47.73 35.90 0.2239 1.0020 1.0020